OpenStudy (anonymous):

Write a point-slope equation for the line that passes through the point (1, 6) and is parallel to the line given by y = 5x + 25

OpenStudy (anonymous):

Gradient of the line = Gradinet of the given line = 5 (because they are parallel) Equation of the line: (y-y1) = m (x-x1), where m is gardient of line =5, and x1y1 is a point on the line. (y-6) = 5 (x-1)

OpenStudy (anonymous):

Ok so u know: y-y1=m(x-x1) Tell me what m, x1 & y1 are :)

OpenStudy (anonymous):

x is 5 and y is 25 are y is -6 and xis-1 and m not sure

OpenStudy (anonymous):

i think u got a little mixed up there :p ok so first when it says Parallel Slope, that means it's SAME slope always! :) So u would just take the slope out of y=5x+25, which is 5 and use it in the point-slope formula also (x1,y1) is your point so in (1,6) x1=1, y1=6 Now can u put it point-slope formula?

OpenStudy (anonymous):

so y-5=1(x+6)

OpenStudy (anonymous):

close, ok so u know: M=5 X1=1 Y1=6 y-y1=m(x-x1) Think u can get it right this time? :)

OpenStudy (anonymous):

y-5=6(x=1)

OpenStudy (anonymous):

Close